OpenStudy (anonymous):

find the product of (2x-5)(x^2-4x+3)

OpenStudy (anonymous):

(2x-5)(x^2-4x+3) =2x^3-8x^2+6x-5x^2+20x-15 =2x^3-8x^2-5x^2+6x+20x-15 =2x^3-3x^2+26x-15

OpenStudy (anonymous):

Just multiply each term in the first parenthesis with each term in the second parenthesis. Then simplify. (As show above by jedai17)